4. Let Y and Z be independent uniform random variables on the interval [0,1]. Let X Z (a) Compute E(XTY). (b) Compute E(X).
Answer #1

a)

E(X|Y)

f(x|y) = f(x,y) /f(y) = f(x)   as X and Y are independent

уч!

b)

E(X) = E(E(X|Y)) = E(1/(y+1)) =

dy In (2 Decima 0.69315...)